### v1

@@ 4.1 Simulation
In this subsection, we provide two sets of simulations to assess the size and power properties of our sharp test under various DGPs in finite samples. Throughout this section, we ran $1000$ replications for each simulation design, and the bootstrap sample size is chosen to be $B=800$. We set $a_{n}=0.15\ln n$ and $B_{n}={0.85\ln n}/{\ln\ln n}$, as in Hsu, Liu, and Shi (2019). We choose $Q_{P}=5$ and $Q_{Y}=5$ (for continuous $Y$) or $Q_{Y}=2$ (for binary $Y$). We set the infinitesimal constant $\eta=10^{-6}$ and the constant $\epsilon=10^{-6}$ (see the definition of $\hat{\sigma}^{2}_{d,\epsilon}(\ell)$ in 3.1-4). The simulation results are not sensitive to these constants.
@@ 4.1.1 Binary outcome
The first set of simulations is based on a DGP introduced in FLL (online appendix, page 22). In this set of simulations, we mimic the random assignment of $n$ defendants to a pool of $J$ judges, ensuring an equitable distribution of $\frac{n}{J}$ defendants to each judge. As in FLL, the severity probability of each judge $j$ is set as follows:
$p_{j}=p_{a}+\frac{j-1}{J-1}(1-p_{a}-p_{n})$
Here, $p_{a}$ and $p_{n}$ stand for the fraction of always and never treated defendants, respectively. FLL consider a binary outcome model where the outcome $Y\in\{0,1\}$ satisfies the following condition:
$\mathbb{E}\left[Y\mid p_{j}\right]=\frac{1-(1-\lambda)(p_{n}+p_{a})}{1-(p_{n}+p_{a})}p_{j}-\frac{\lambda}{1-(p_{n}+p_{a})}p_{a}.$
The parameter $\lambda$ dictates the extent of deviation from the exclusion restriction assumption. When $\lambda=0$, there is no violation of the judge leniency design assumptions. Consequently, for $\lambda=0$, the simulations aim at assessing the size property of the two different tests. On the other hand, $\lambda>0$ signifies a departure from the exclusion restriction assumption, with higher (absolute) values indicating a more pronounced deviation. Like in the original paper, we adopt the following parametrization for the fraction of always and never treated $p_{n}=p_{a}=0.2$. Meanwhile, we vary the value of $\lambda$ within the range of 0 to 1.
(a) $-\mathbb{E}[Y(1-D)|P=p]$
@@
[…]
@@
FLL discuss how one can improve the power of their testing methodology by considering more stringent upper bounds on the largest possible treatment effects (i.e. using a smaller value of $K$). For instance, in their empirical application of a binary outcome model–where the maximum treatment effect is set at 1–they advocate exploring smaller permissible maximum treatment effect values. However, if $K$ is set to be too small, then FLL’s test can have server size distortion. Indeed, Figure 2-(b) graphically represents this situation by plotting the rejection rate associated with FLL’s nonparametric test under two additional cases: when the maximum allowable treatment effect $K$ is set at 0.8 and 0.4, respectively. The striking observation is the conclusions drawn from these scenarios can be misleading, as they suggest an excessive over-rejection of the assumptions even when those assumptions are indeed satisfied. For example, if one sets $K=0.4$, then the rejection rate is always $100\%$ whenever the sample size is greater or equal to $1000$. As a matter of fact, the rejection we observe from Figure 2-(b) reflects that the ad-hoc imposed magnitude of the treatment effect is not correct, but the underlying exclusion restriction holds. Our test is immune to this problem since it does not require pre-specifying the magnitude order of the unknown treatment effect.
@@ 4.1.2 Continuous outcome
The second set of simulations aims to show the performance of our test in detecting violations of the judge leniency design when the outcome is continuous and unbounded. Let $(U_{0},U_{1},U,Z^{*})\sim N(\boldsymbol{\mu},\Sigma)$, where $\boldsymbol{\mu}=(\mu_{0},\mu_{1},\mu_{U},\mu_{Z})^{\prime}$ is a vector of means, and $\Sigma$ is a covariance matrix. For generic random variables $A$ and $B$, let $\sigma_{A}^{2}$ be the variance of $A$ and $\rho_{A,B}$ be the correlation coefficient between $A$ and $B$. In this design, we set $\sigma_{A}=1$ for all $A\in\{U_{1},U_{0},U,Z^{*}\}$. We let $\rho_{U_{0},U}=-0.5$, $\rho_{U_{1},U}=0.5$, $\rho_{U,Z}=0$, $\rho_{U_{1},U_{0}}=0$, $\rho_{U_{1},Z}=\delta_{1}$, and $\rho_{U_{0},Z}=\delta_{1}$. To create discrete judges or IV $Z$, we set
$Z=F_{Z^{*}}^{-1}\left(\min\left\{\ell=1,2,\cdots,L-1:\left|F_{Z^{*}}(Z^{*})-\frac{\ell}{L}\right|\right\}\right),$
@@
    which indicates that there are two groups of judges, with each group having distinct skills (or preferences) in assigning treatment. This is in clear violation of the monotonicity assumption, which requires all judges to have the same skill (Chan, Gentzkow, and Yu, 2022).
@@
Figure 4, on the other hand, plots the testable implication used in FLL. The left side panels plot $\mathbb{E}[Y|P=p]$ for each of the $p\in\{p_{1},p_{2},\cdots,p_{20}\}$ (sorted in increasing order) for each of the four designs. The right panels plot the “numerical derivative” of the form $\frac{\mathbb{E}[Y|P=p_{j}]-\mathbb{E}[Y|P=p_{j-1}]}{p_{j}-p_{j-1}}$ against $\{p_{2},\cdots,p_{20}\}$. The FLL testable implications require that the curves in the right-hand side panels be bounded between $[-K,K]$, where $K$ again is the difference between the upper and lower bounds of the support. Note that in this example, the outcomes have unbounded support and, therefore, $K=+\infty$. If we choose $K$ as a large number, then it is apparent all four designs satisfy FLL’s testable implication. Hence, we expect no rejection for designs 2-4 albeit they violated the identifying assumptions unless $K$ is set to be relatively small.
We proceed by implementing our sharp test and FLL’s nonparametric test. This comparison is conducted across various parameter values and sample sizes. Specifically, we consider a size design (Size $\delta_{1}=\delta_{2}=\delta_{3}=0$), violation of independence (Power1 $\delta_{1}=-0.5,\delta_{2}=\delta_{3}=0$), violation of monotonicity (Power2 $\delta_{2}\neq 0,\delta_{1}=\delta_{3}=0$), and violation of exclusion (Power3 $\delta_{3}=-0.5,\delta_{1}=\delta_{2}=0$). For each violation, we consider situations with covariates ($\beta_{1}=\beta_{0}=1$) or without covariates ($\beta_{1}=\beta_{0}=0$ ). To implement FLL’s test, we set $K$ to be the difference between sample maximum ($y_{max}$) and minimum $(y_{min})$: $\Delta_{y}\equiv y_{max}-y_{min}$. We also consider $K=\frac{\Delta_{y}}{8}$ and $K=\frac{\Delta_{y}}{16}$. The results are summarized in LABEL:table:_types_of_violations_no_x.
@@
Regarding the size property, all tests control the size except the FLL test when $K$ is set to be very small. Our test and the FLL test with $K=\Delta_{y}$ and $K=\frac{\Delta_{y}}{8}$ are conservative. When one sets $K=\frac{\Delta_{y}}{16}$, the rejection probability of FLL’s test increases quickly even when all the assumptions are satisfied (the first design). This is unsurprising because a very small $K$ essentially introduced another severe misspecification to the model. However, when examining the power property of the three tests, we see clearly that our test outperforms FLL’s tests by a large margin. The proposed sharp test has enough power to detect the violation of any of the three assumptions (independence, exclusion and monotonicity). In particular, the rejection rates for our sharp test quickly increase with sample size, surpassing 90% for all cases when the sample size reaches $2000$ (or $100$ cases per judge). Note that in this simulation, the parametric form of the propensity score is correctly specified (except for Power2 when monotonicity is violated); hence, the high power of our test is not because of misspecification of $P(z,\theta_{0})$. In contrast, FLL’s test has low power performance unless we set $K$ as a small value, which, on the other hand, induces size distortion.
@@ 4.2 Empirical illustration
In the Philadelphia court system, following an arrest, individuals are taken to one of seven city police stations for a videoconference interview by Pretrial Services, assessing risk factors and financial details for public defence eligibility. Utilizing this information, Pretrial Services assigns arrestees to a bail recommendation grid. Bail hearings, conducted by magistrates every four hours over videoconference, involve a brief process where charges are explained, next court appearances are specified, eligibility for a court-appointed defence attorney is determined, and bail amounts are set based on arrest details, interviews, criminal history, guidelines, and input from representatives. Magistrates hold broad authority to assign bail, which can fall into categories such as release without payment, cash bail with a 10% deposit, or no bail at all.
Stevenson (2018)’s research design leverages the varying magistrate tendencies to assign affordable bail as an instrument to study detention’s impact on cases’ outcomes. To answer the research questions, the author uses data from the court records of the Pennsylvania Unified Judicial System, obtained through web-scraping of public records in PDF format, which are then transformed for statistical analysis. The dataset encompasses arrests in Philadelphia, where charges were filed between September 13, 2006, and February 18, 2013. The final dataset includes 331,971 cases and eight randomly assigned judges, with each observation pertaining to a specific criminal case.
In what follows, we focus on the aggregate dataset (all criminal cases together) and four primary categories of criminal cases in the data: aggressive assault, robbery, drug sale, and drug possession. These four criminal cases we consider in isolation constitute 43% of the total cases. In Figure 5, we present a scatter plot with a fitted polynomial to illustrate whether the anticipated implications of the judge leniency design framework are satisfied for the considered categories of criminal cases. The graphs indicate $\mathbb{E}[YD|P=p]$ and $\mathbb{E}[-Y(1-D)|P=p]$ are most likely to be non-decreasing for the aggressive assault case. [^14] Note all the outcome variables are binary. Therefore, the close interval we use for the Theorem 2.1 is $1\{0<Y\leq 1\}$, which equals to $Y$. The non-decreasing shape of the functions is unclear for the other types of criminal categories. Although this graphical representation does not constitute a formal test, it offers an intuitive insight. Specifically, it suggests the assumptions are the least likely to be violated for the aggressive assault case, while the drug possession case shows the highest likelihood of violations of the assumptions of the judge leniency design.
@@
We observe a relatively large set of covariates, including fixed effects for year, month, and day of the week. We, consequently, implement the semi-parametric version of our test. For comparison, we also implement FLL’s nonparametric and semi-parametric tests. The results of the three tests are presented in LABEL:table:_empirical_application for both the aggregate dataset and separately for each of the four crime categories aforementioned. The nonparametric and semi-parametric tests introduced by FLL indicate the validity of judge leniency design cannot be rejected either conditioning on each crime category or the aggregate data set at 10% level, despite that the shape of $\mathbb{E}[YD|P=p]$ and $\mathbb{E}[-Y(1-D)|P=p]$ for the drug possession type suggests the opposite. In contrast, our novel test yields results that align with expectations. For instance, the sharp test does not indicate a rejection of the validity of the judge leniency design assumptions for the aggressive assault. However, for all three other types of offences, our test rejects the validity of the judge leniency design, which means using the Wald estimand or the MTE approach for those cases will lead to inconsistent estimates. Finally, we see no evidence to refute the assumptions underpinning the judge leniency design when applying our sharp test to the aggregate data set. This outcome may be influenced by the notably high proportion of aggressive assault cases within the dataset compared to other categories. Our result also ascertains that the exclusion restriction or monotonicity can hold for some crime categories but not others, suggesting controlling the crime type is important in practice.
@@
All 0.821 0.056 1.000
Aggressive assault 0.913 0.996 1.000
Robbery 0.033 1.000 0.898
Drug sale 0.005 0.116 1.000
Drug possession 0.000 0.929 1.000
Notes: This table reports the results of the statistical tests using Stevenson (2018)’data, including time fixed effects as controls. Specifically, the considered controls are year, month, and day of the week fixed effects. Sharp Test stands for our novel semi-parametric test developed in this paper, while FLL-Nonp and FLL-Semip represent the nonparametric and semi-parametric tests of FLL, respectively. We follow FLL’s empirical specification and use 5 knots for their semi-parametric test.
5 Partial Exclusion and Monotonicity
The rejection of the sharp test means the judge’s leniency design assumptions are too stringent for the data. In these cases, it may be desirable to relax some of these assumptions to salvage the model. In this section, we consider partial exclusion and monotonicity to weaken the assumptions. Consider a general form of the potential outcome model:
$Y$ $=$ $\tilde{Y}_{1}D+\tilde{Y}_{0}(1-D),$
$\tilde{Y}_{d}$ $=$ $\sum_{z\in\mathcal{Z}}Y_{d}(z)1\{Z=z\},$
$D$ $=$ $\sum_{z\in\mathcal{Z}}D_{z}1\{Z=z\}.$
Let us decompose $Z$ into two components $Z\equiv(Z_{I},Z_{c})$ where $Z_{I}$ is the judge fixed effect, and $Z_{c}$ is a vector of judge observable characteristics, such as experience, race, and political party, among others.

### v2

@@ 4.1 Simulation
In this subsection, we provide two sets of simulations to assess the size and power properties of our sharp test under various DGPs in finite samples. Throughout this section, we ran $1000$ replications for each simulation design, and the bootstrap sample size is chosen to be $B=800$. We set $a_{n}=0.15\ln n$ and $B_{n}={0.85\ln n}/{\ln\ln n}$, as in Hsu, Liu, and Shi (2019). We choose $Q_{P}=5$ and $Q_{Y}=5$ (for continuous $Y$) or $Q_{Y}=2$ (for binary $Y$). We set the infinitesimal constant $\eta=10^{-6}$ and the constant $\epsilon=10^{-6}$ (see the definition of $\hat{\sigma}^{2}_{d,\epsilon}(\ell)$ in 3.1-4).
@@ 4.1.1 Binary outcome
The first set of simulations is based on a DGP introduced in FLL (online appendix, page 22). In this set of simulations, we mimic the random assignment of $n$ defendants to a pool of $J$ judges, ensuring an equitable distribution of $\frac{n}{J}$ defendants to each judge. As in FLL, the severity probability of each judge $j$ is set as follows: p_j = p_a + j-1J-1(1-p_a - p_n) Here, $p_{a}$ and $p_{n}$ stand for the fraction of always and never treated defendants, respectively. FLL consider a binary outcome model where the outcome $Y\in\{0,1\}$ satisfies the following condition: E [ Y ∣p_j ] = 1 - (1-λ)(pn+ pa)1 - (pn+ pa) p_j - λ1-(pn+ pa) p_a. The parameter $\lambda$ dictates the extent of deviation from the exclusion restriction assumption. When $\lambda=0$, there is no violation of the judge leniency design assumptions. Consequently, for $\lambda=0$, the simulations aim at assessing the size property of the two different tests. On the other hand, $\lambda>0$ signifies a departure from the judge leniency design assumption, with higher (absolute) values indicating a more pronounced deviation. Like in the original paper, we adopt the parametrization for the fraction of always and never treated $p_{n}=p_{a}=0.2$. Meanwhile, we vary the value of $\lambda$ within the range of 0 to 1. Note that the parameter $\lambda$ directly governs the shape of the function $\mathbb{E}[Y|P=p]$. The nonzero value of $\lambda$ can potentially be generated by violations of one of the three assumptions (or their combinations).
(a) $\mathbb{E}[Y(1-D)|P=p]$
@@
In Figure 2-(a), we report the size property for our sharp test and FFL’s test at 5% significance level (when $\lambda=0$). The simulation designs involve twenty judges and varying sample sizes, ranging from 500 defendants (equivalent to 50 defendants per judge) to 5500 defendants (equivalent to 550 defendants per judge). The plot reveals that both tests control size well in the aforementioned DGP. Specifically, it is evident from the graph that the rejection rate of our sharp test is controlled by and close to the nominal level of 5%. Conversely, the nonparametric test proposed in FLL consistently yields rejection rates close to zero when setting the tuning parameter K=1K=1.151515Recall the outcome variable is binary; hence, the largest possible absolute value for the treatment effect is $1$. These results correspond to Figures 9 and 10 in the online appendix of FLL, where the rejection probabilities are nearly zero for various sample sizes when $\lambda=0$.
@@
FLL discuss how one can improve the power of their testing methodology by considering more stringent upper bounds on the largest possible treatment effects (i.e., using a smaller value of $K$). For instance, in their empirical application of a binary outcome model–where the maximum treatment effect is set at 1–they advocate exploring smaller permissible maximum treatment effect values. However, if $K$ is set to be too small, then FLL’s test can have server size distortion. Indeed, Figure 2-(b) graphically represents this situation by plotting the rejection rate associated with FLL’s nonparametric test under two additional cases: when the maximum allowable treatment effect $K$ is set at 0.8 and 0.4, respectively. The striking observation is that the conclusions drawn from these scenarios can be misleading, as they suggest an excessive over-rejection of the assumptions even when those assumptions are indeed satisfied. For example, if one sets $K=0.4$, then the rejection rate is always $100\%$ whenever the sample size is greater or equal to $1000$. As a matter of fact, the rejection we observe from Figure 2-(b) reflects that the ad-hoc imposed magnitude of the treatment effect is not correct, but the underlying exclusion restriction holds. Our test is immune to this problem since it does not require pre-specifying the magnitude order of the unknown treatment effect.
@@ 4.1.2 Continuous outcome
The second set of simulations aims to show the performance of our test in detecting violations of the judge leniency design when the outcome is continuous and unbounded. Let $(U_{0},U_{1},U,Z^{*})\sim N(\boldsymbol{\mu},\Sigma)$, where $\boldsymbol{\mu}=(\mu_{0},\mu_{1},\mu_{U},\mu_{Z})^{\prime}$ is a vector of means, and $\Sigma$ is a covariance matrix. For generic random variables $A$ and $B$, let $\sigma_{A}^{2}$ be the variance of $A$ and $\rho_{A,B}$ be the correlation coefficient between $A$ and $B$. In this design, we set $\sigma_{A}=1$ for all $A\in\{U_{1},U_{0},U,Z^{*}\}$. We let $\rho_{U_{0},U}=-0.5$, $\rho_{U_{1},U}=0.5$, $\rho_{U,Z}=0$, $\rho_{U_{1},U_{0}}=0$, $\rho_{U_{1},Z}=\delta_{1}$, and $\rho_{U_{0},Z}=\delta_{1}$. To create discrete judges or IV, we set
$Z=F_{Z^{*}}^{-1}\left(\frac{\ell(Z^{*})}{L}\right),\quad\ell(Z^{*})=\mathop{\mathrm{argmin}}_{\ell=1,2,\cdots,L-1}\left|F_{Z^{*}}(Z^{*})-\frac{\ell}{L}\right|.$
@@
    which indicates that there are two groups of judges, each with distinct skills (or preferences) in assigning treatment. This is in clear violation of the monotonicity assumption, which requires all judges to have the same skill (Chan, Gentzkow, and Yu, 2022).
@@
Figure 8, on the other hand, plots the testable implication used in FLL. The left side panels plot $\mathbb{E}[Y|P=p]$ for each of the $p\in\{p_{1},p_{2},\cdots,p_{20}\}$ (sorted in increasing order) for each of the four designs. The right panels plot the “numerical derivative” of the form $\frac{\mathbb{E}[Y|P=p_{j}]-\mathbb{E}[Y|P=p_{j-1}]}{p_{j}-p_{j-1}}$ against $\{p_{2},\cdots,p_{20}\}$. The FLL testable implications require that the curves in the right-hand side panels be bounded between [K,K][-K,K], where $K$ again is the difference between the upper and lower bounds of the support. Note that in this example, the outcomes have unbounded support and, therefore, $K=+\infty$. If we choose $K$ as a large number, then it is apparent that all four designs satisfy FLL’s testable implication. Hence, we expect no rejection for designs 2-4, albeit they violate the identifying assumptions unless $K$ is set to be relatively small.
@@
We proceed by implementing our sharp test and FLL’s nonparametric test. This comparison is conducted across various parameter values and sample sizes. Specifically, we consider a size design (Size $\delta_{1}=\delta_{2}=\delta_{3}=0$), violation of independence (Power1 $\delta_{1}=-0.5,\delta_{2}=\delta_{3}=0$), violation of monotonicity (Power2 $\delta_{2}\neq 0,\delta_{1}=\delta_{3}=0$), and violation of exclusion (Power3 $\delta_{3}=-0.5,\delta_{1}=\delta_{2}=0$). For each violation, we consider situations with covariates ($\beta_{1}=\beta_{0}=1$) or without covariates ($\beta_{1}=\beta_{0}=0$ ). When there are covariates, we use Carr and Kitagawa (2021)’s method to control for covariates, as discussed in the previous section. To implement FLL’s test, we set $K$ to be the difference between sample maximum ($y_{max}$) and minimum $(y_{min})$: $\Delta_{y}\equiv y_{max}-y_{min}$. We also consider $K=\frac{\Delta_{y}}{8}$ and $K=\frac{\Delta_{y}}{16}$. The results are summarized in LABEL:table:_types_of_violations_no_x.
@@
Regarding the size property, all tests control the size except the FLL test when $K$ is set to be very small. Our test and the FLL test with $K=\Delta_{y}$ and $K=\frac{\Delta_{y}}{8}$ are conservative. When one sets $K=\frac{\Delta_{y}}{16}$, the rejection probability of FLL’s test increases quickly even when all the assumptions are satisfied (the first design). This is unsurprising because a very small $K$ essentially introduced another severe misspecification to the model. However, when examining the power property of the three tests, we see clearly that our test outperforms FLL’s tests by a large margin. The proposed sharp test has enough power to detect the violation of any of the three assumptions (independence, exclusion, and monotonicity). In particular, the rejection rates for our sharp test quickly increase with sample size, surpassing 90% for all cases when the sample size reaches $2000$ (or $100$ cases per judge). Note that in this simulation, the parametric form of the propensity score is correctly specified (except for Power2 when monotonicity is violated); hence, the high power of our test is not because of misspecification of $P(z,\theta_{0})$. In contrast, FLL’s test has low power performance unless we set $K$ as a small value, which, on the other hand, induces size distortion.
@@ 4.2 Empirical illustration
In the Philadelphia court system, following an arrest, individuals are taken to one of seven city police stations for a video conference interview by Pretrial Services, which assesses risk factors and financial details for public defense eligibility. Utilizing this information, Pretrial Services assigns arrestees to a bail recommendation grid. Bail hearings, conducted by magistrates every four hours via video conference, involve a brief process where charges are explained, next court appearances are specified, eligibility for a court-appointed defense attorney is determined, and bail amounts are set based on arrest details, interviews, criminal history, guidelines, and input from representatives. Magistrates hold broad authority to assign bail, which can fall into categories such as release without payment, cash bail with a 10% deposit, or no bail at all.
Stevenson (2018)’s research design leverages the varying magistrate tendencies to assign affordable bail as an instrument to study detention’s impact on case outcomes. To answer the research questions, the author utilizes data from the court records of the Pennsylvania Unified Judicial System, obtained through web scraping of public records in PDF format, which are then transformed for statistical analysis. The dataset encompasses arrests in Philadelphia, where charges were filed between September 13, 2006, and February 18, 2013. The final dataset includes 331,971 cases and eight randomly assigned judges, with each observation pertaining to a specific criminal case. As noted in Stevenson (2018), the shift-rotation system at the Philadelphia court forms the basis for such randomness.
In what follows, we focus on the aggregate dataset (all criminal cases together) and four primary categories of criminal cases in the data: aggressive assault, robbery, drug sale, and drug possession. These four criminal cases we consider in isolation constitute 43% of the total cases. In Figure 5, we present two scatter plots for each crime category: $\{(p_{j},\mathbb{E}[YD|P=p_{j}])\}_{j=1}^{8}$ and $\{(p_{j},\mathbb{-}E[Y(1-D)|P=p_{j}])\}_{j=1}^{8}$, along with a fitted polynomial to illustrate whether the anticipated implications of the judge leniency design framework are satisfied for the considered categories of criminal cases. The graphs indicate $\mathbb{E}[YD|P=p]$ and $\mathbb{E}[-Y(1-D)|P=p]$ are most likely to be non-decreasing for the aggressive assault case. [^16] Note all the outcome variables are binary. Therefore, the close interval we use for the Theorem 2.1 is $1\{0<Y\leq 1\}$, which equals to $Y$. The non-decreasing shape of the functions is unclear for the other types of criminal categories. Although this graphical representation does not constitute a formal test, it offers an intuitive insight. Specifically, it suggests that the assumptions are the least likely to be violated in the aggressive assault case, while the drug possession case shows the highest likelihood of violating the assumptions of the judge leniency design.
@@
We observe a relatively large set of covariates, including fixed effects for year, month, and day of the week. We, consequently, implement the semi-parametric version of our test. For comparison, we also implement FLL’s nonparametric and semi-parametric tests. The results of the three tests are presented in LABEL:table:_empirical_application for both the aggregate dataset and separately for each of the four crime categories aforementioned. The nonparametric test introduced by FLL indicates the validity of judge leniency design cannot be rejected either conditioning on each crime category or the aggregate data set at 10% level, despite that the shape of $\mathbb{E}[YD|P=p]$ and $\mathbb{E}[-Y(1-D)|P=p]$ for the drug possession type suggests the opposite. In contrast, our novel test yields results that align with expectations. For instance, the sharp test does not indicate a rejection of the validity of the judge leniency design assumptions for the aggressive assault. However, for all three other types of offenses, our test rejects the validity of the judge leniency design. Meanwhile, FLL’s semi-parametric test rejects the category of aggregate assault. [^17] For FLL’s semi-parametric test, we fit the regression function $\mathbb{E}[Y|P=p]$ by B-spline with three knots. The results for other numbers of knots are reported in the appendix. The reported p-value is the “combined p-value” of the fit component and slope component of the test, and we can see from LABEL:table:_FLL_p_values in the online appendix that the rejection is mostly generated by the fit component. These results suggest that using the Wald estimand or the MTE approach for those cases will lead to inconsistent estimates of the causal effects of interest.
Finally, we see no evidence to refute the assumptions underpinning the judge leniency design when applying our sharp test to the aggregate dataset. This outcome may be influenced by the notably high proportion of aggressive assault cases within the dataset compared to other categories. Our result also ascertains that the exclusion restriction or monotonicity can hold for some crime categories but not others, suggesting that controlling the crime type is important in practice.
@@
All 0.821 0.056 0.114
Aggressive assault 0.913 0.996 0.015
Robbery 0.033 1.000 0.109
Drug sale 0.005 0.116 0.180
Drug possession 0.000 0.929 0.610
Notes: This table reports the results of the statistical tests using Stevenson (2018)’data, including time fixed effects as controls. Specifically, the considered controls are year, month, and day-of-the-week fixed effects. Sharp Test stands for our novel semi-parametric test developed in this paper, while FLL-Nonp and FLL-Semip represent the nonparametric and semi-parametric tests of FLL (three knots B-spline), respectively.
5 Salvage the Model under Weaker Assumptions
The rejection of the sharp test means that the judge’s leniency design assumptions are too stringent for the data. In this case, relaxing some of these assumptions is required to salvage the model. There are different ways to relax a model’s assumptions. One way is to maintain the same estimand used in the stringent model and ask under what conditions this estimand can still be interpreted causally. The model relaxation recently entertained by FLL falls into this second approach, providing alternative conditions under which the 2SLS could still have a causal interpretation when Assumptions 2.1, 2.2 and 2.3 are too stringent for the data. In Section 5.1, we revisit the average exclusion assumption proposed by FLL and show it is a special case of a zero-covariance condition: a restriction that may not always be justifiable in all empirical settings.
There is another approach that focuses on a well-defined policy-relevant parameter and examines how this parameter could be point-identified or set-identified using weaker and more credible assumptions. In such a case, the parameter of interest remains the same, but the (set) estimands may vary depending on the credible assumptions one would be willing to maintain. We will discuss this approach in Section 5.2.181818In this section, we mainly focus on the case in which the exclusion or monotonicity assumption is violated. When the random assignment assumption is violated, one can consider a partial identification, see Mourifié and Wan (2025).
5.1 Average Exclusion and Monotonicity
We first revisit the average exclusion and monotonicity conditions. For simplicity, suppose $Z$ has finite support as in FLL such that $\mathcal{Z}=\{1,2,\cdots,J\}$. The general form of the potential outcome model is,
$Y=\tilde{Y}_{1}D+\tilde{Y}_{0}(1-D),\quad\tilde{Y}_{d}=\sum_{z\in\mathcal{Z}}Y_{d}(z)1\{Z=z\},\quad D=\sum_{z\in\mathcal{Z}}D_{z}1\{Z=z\}.$
FLL proposes to relax Assumption 2.2 and Assumption 2.3 with the average exclusion restriction and the average monotonicity assumption, respectively: